# Identification of graphite with perfect rhombohedral stacking by electronic Raman scattering

### v2

@@ Results
Although the 2D peak shape converges with large , its integrated intensity ratio should be used as a first tool to rapidly screen graphite flakes for RG containing flakes. Furthermore, determining the precise number of graphene layers in the flake is a crucial step in identifying RG without stacking faults, therefore we describe this step in detail below. The layer number can be easily determined by optical contrast 32; 33, in the case of <10 layers. For thicker samples atomic force microscopy (AFM) can be used, provided care is taken during the measurement, to counter any anomalous height signal stemming from the difference in properties of the graphite and substrate 29. Fig. 2a-c illustrates the thickness measurement of a graphite flake exfoliated onto a Si / SiO substrate, where a graphene layer slightly extends near the top in the optical microscopy image (Fig. 2a). For accurate height measurement using dynamic atomic force microscopy (AFM) methods, it is essential to measure the flake’s thickness relative to this bottom graphene layer, or in an area where the flake folds back onto itself 29. Raman spectroscopy confirms that the protruding graphene is indeed a single layer. The well-known height anomaly in dynamic (tapping mode) AFM images 29 is evident in Fig. 2b. Here, the graphene layer’s height, relative to the SiO substrate, registers at 1.15 nm instead of the expected 0.33 nm. Conversely, the single-layer step in the middle of the flake is accurately measured, aligning with the van der Waals distance of graphite, as illustrated in Fig. 2c. Therefore, when measured relative to the bottom graphene layer, the flake’s thickness is determined to be 2.32 nm, corresponding to seven graphene layers. Including the bottom layer, our flake comprises regions with eight and seven layers, respectively. The layer number is also reproduced from optical contrast measurements of the layer thickness.
The AFM topography image in Fig. 2c reveals a domain wall within the flake. A Raman spectroscopy map measured across this flake (Fig. 2d) shows that the domain wall separates areas with wide and narrow 2D peaks. In both the 7 and 8 layer regions, the larger 2D peak intensity ratio suggest rhombohedral stacking. Fig. 2e features averaged Raman spectra from the hexagonal, and the 7 and 8 layer rhombohedral regions. The bottom panel focuses on the Raman spectra’s background signal. In addition to the well known low-intensity phonon modes (M, iTALO, iTOTA, LOLA), 34; 35; 22 the background reveals a wide peak on both the 7 and 8 layer rhombohedral areas. The halfwidths of these peaks is in the 500 cm regime, consistent with the temperature broadening at room temperature ( meV for K, where is the Boltzmann constant). Their absence in the hexagonal region implies that their origin is electronic Raman scattering, because the ERS in hexagonal graphite is mostly flat 27; 28 due to a lack of a bulk gap and other sharp DOS features.
@@
A polarization dependent measurement can also be used to verify that the broad peaks observed in our sample originate from ERS. If the incoming excitation laser is linearly polarized, the optical selection rules of the ERS process result in a perpendicular linear polarization of the scattered photons 36; 31; 27. Fig. 3b displays the 2D peak alongside the broad ERS response for the 7 and 8 layer regions of our flake. Unlike the spectrum in Fig. 2e, a polarizer is now positioned in the path of the scattered light, oriented perpendicularly to the polarization of the incoming excitation laser. This "crossed polarizer" setup amplifies the relative visibility of the ERS signal, which becomes 10% of the 2D peak intensity, as shown in Fig. 3b. Consequently, we can confirm that the broad peaks stem from ERS by comparing measurements from the same location under both crossed and parallel polarizer configurations, as depicted in Fig. 3c. In the parallel configuration, the broad peaks are absent, whereas they are clearly present in the crossed polarizer setup. This comparison can also be used to separate the ERS peaks from the phonon peaks 34; 35. By performing the same measurement in the crossed and parallel configuration we can subtract the two curves, which leaves us with the ERS signal. We attribute the peak exhibiting the lower Raman shift to the transition, and the peak at higher shift is identified as the process. To extract the ERS signal, we first subtract a linear background and normalize to a phonon peak, in this example (Fig. 3c) to the 2D + G peak at 4300 cm. After this we subtract the two spectra, which leaves us with the ERS signal. We fit the resulting ERS peaks by Gaussians to determine their Raman shift. The result of this procedure is shown in Fig. 3c. A detailed description of the fitting process can be found in section S6 of the Supplementary Information and the shared data 37.
We apply this procedure to samples with thicknesses varying from 3 to 12 graphene layers. Each sample undergoes meticulous examination using atomic force microscopy (AFM) to ensure accurate layer count. Additionally, we identify few layer samples through optical contrast measurements 32; 33. The ERS peaks of these samples can be seen in Fig. 4a. We can observe a clear trend: with increasing layer number the and peaks continuously shift to lower energy (lower Raman shift). By plotting these peak positions extracted by Gaussian fitting versus the layer number this trend is even more evident (see Fig. 4b). In certain cases the ERS peak positions have a larger estimated error (120 cm) due to lower signal to noise or due to the overlap of the ERS peak with the 2D mode in the case of 9 and 10 layers. These measurements align within our error margins with the calculated positions of ERS peaks from references: 27; 28 (see Fig. 4b). The measured ERS peak positions are also supplied as a table in the Methods (see Table 1).
@@
Next we check the homogeneity of the ERS peak position over the flake. Taking the difference in the crossed and parallel polarization measurements across the crystal, we plot the Raman intensity of the transition for the 8 and 7 layer areas. It becomes clear that the largest intensity of the ERS peak is located exactly on the rhombohedral 7 and 8 layer regions respectively (Fig. 5a). To check the homogeneity of the rhombohedral regions, in Fig. 5b we plot the position of the ERS peak. The standard deviation within both regions is under 5 cm. If a stacking fault were present in these regions, a change in peak energy two orders of magnitude greater would be expected. This is because the shift in ERS peak position, due to a stacking fault, is of the order of the change in energies with adding another layer to perfectly stacked RG 27. Thus, the change in peak position would be least 300 cm in the presence of a stacking fault.
@@ Methods
We exfoliate samples using "blue tape" (Ultron systems P/N: 1008R-8.0), but other tapes work equally well. As substrate we used Si wafers (90 nm SiO). Graphite samples were purchased from NGS Trading & Consulting GmbH (www.graphit.de). Based on hundreds of flakes investigated by Raman measurements, roughly 40% of them have some rhombohedral domains. We performed AFM measurements, using an NX10 microscope from Park Systems in tapping mode (non-contact mode).
@@
For Raman measurements we use a Witec 300rsa+ confocal Raman system, using 488, 532, and 633 nm laser excitations. All data shown in the main text are measured using a diffraction grating with 600 lines mm. Raman measurements are first analysed using the Witec data processing software supplied with the confocal Raman system. We performed final data processing and generated the figures, using the open-source Python tool: Ramantools 43.
| Layer Number | ERS peak position (cm) | ERS peak position (cm) |
@@
STM measurements are carried out at a temperature of 9 K, using an instrument from RHK (PanScan Freedom), with a base pressure of 510 Torr. Tunnelling conductance measurements were performed, using a Lock-in amplifier, at a frequency of 1372 Hz and a bias modulation amplitude of 5.5 mV. Samples investigated by STM are exfoliated flakes, supported on a Si/SiO substrate and contacted using In spikes 44. Data analysis of the STM measurements is carried out using the open-source Python tool: RHKPY 45.
@@
The optimized geometry and ground state Hamiltonian and overlap matrix elements of each structure were self consistently obtained by the SIESTA implementation of density functional theory (DFT) 46; 47; 48. SIESTA employs norm-conserving pseudopotentials to account for the core electrons and linear combination of atomic orbitals to construct the valence states. For all cases the considered samples were separated with a minimum of 1.35 nm thick vacuum in the perpendicular direction. The generalized gradient approximation of the exchange and the correlation functional was used with Perdew-Burke-Ernzerhof parametrization 49 with a double- polarized basis set. The geometry optimizations were performed until the forces were smaller than 0.1 eV nm . The geometry of the considered structures were optimized for every configuration, initiated from the experimental in-plane lattice constant = 0.246 nm and out-of-plane lattice constant = 0.670 nm of hexagonal graphite. During the geometry relaxation the real-space grid was defined with an equivalent energy cutoff of 400 Ry and the Brillouin zone integration was sampled by a 1201201 Monkhorst-Pack -grid 50.

### v3

@@ Results
Although the 2D peak shape converges with large , its integrated intensity ratio should be used as a first tool to rapidly screen graphite flakes for RG containing flakes. Furthermore, determining the precise number of graphene layers in the flake is a crucial step in identifying RG without stacking faults, therefore we describe this step in detail below. The layer number can be easily determined by optical contrast 32; 33, in the case of <10 layers. For thicker samples atomic force microscopy (AFM) can be used, provided care is taken during the measurement, to counter any anomalous height signal stemming from the difference in properties of the graphite and substrate 29. Fig. 2a-c illustrates the thickness measurement of a graphite flake exfoliated onto a Si / SiO2 substrate, where a graphene layer slightly extends near the top in the optical microscopy image (Fig. 2a). For accurate height measurement using dynamic atomic force microscopy (AFM) methods, it is essential to measure the flake’s thickness relative to this bottom graphene layer, or in an area where the flake folds back onto itself 29. Raman spectroscopy confirms that the protruding graphene is indeed a single layer. The well-known height anomaly in dynamic (tapping mode) AFM images 29 is evident in Fig. 2b. Here, the graphene layer’s height, relative to the SiO2 substrate, registers at 1.15 nm instead of the expected 0.33 nm. Conversely, the single-layer step in the middle of the flake is accurately measured, aligning with the van der Waals distance of graphite, as illustrated in Fig. 2c. Therefore, when measured relative to the bottom graphene layer, the flake’s thickness is determined to be 2.32 nm, corresponding to seven graphene layers. Including the bottom layer, our flake comprises regions with eight and seven layers, respectively. The layer number is also reproduced from optical contrast measurements of the layer thickness.
The AFM topography image in Fig. 2c reveals a domain wall within the flake. A Raman spectroscopy map measured across this flake (Fig. 2d) shows that the domain wall separates areas with wide and narrow 2D peaks. In both the 7 and 8 layer regions, the larger 2D peak intensity ratio suggest rhombohedral stacking. Fig. 2e features averaged Raman spectra from the hexagonal, and the 7 and 8 layer rhombohedral regions. The bottom panel focuses on the Raman spectra’s background signal. In addition to the well known low-intensity phonon modes (M, iTALO, iTOTA, LOLA), 34; 35; 22 the background reveals a wide peak on both the 7 and 8 layer rhombohedral areas. The halfwidths of these peaks is in the 500 cm-1 regime, consistent with the temperature broadening at room temperature ( meV for K, where is the Boltzmann constant). Their absence in the hexagonal region implies that their origin is electronic Raman scattering, because the ERS in hexagonal graphite is mostly flat 27; 28 due to a lack of a bulk gap and other sharp DOS features.
@@
A polarization dependent measurement can also be used to verify that the broad peaks observed in our sample originate from ERS. If the incoming excitation laser is linearly polarized, the optical selection rules of the ERS process result in a perpendicular linear polarization of the scattered photons 36; 31; 27. Fig. 3b displays the 2D peak alongside the broad ERS response for the 7 and 8 layer regions of our flake. Unlike the spectrum in Fig. 2e, a polarizer is now positioned in the path of the scattered light, oriented perpendicularly to the polarization of the incoming excitation laser. This "crossed polarizer" setup amplifies the relative visibility of the ERS signal, which becomes 10% of the 2D peak intensity, as shown in Fig. 3b. Consequently, we can confirm that the broad peaks stem from ERS by comparing measurements from the same location under both crossed and parallel polarizer configurations, as depicted in Fig. 3c. In the parallel configuration, the broad peaks are absent, whereas they are clearly present in the crossed polarizer setup. This comparison can also be used to separate the ERS peaks from the phonon peaks 34; 35. By performing the same measurement in the crossed and parallel configuration we can subtract the two curves, which leaves us with the ERS signal. We attribute the peak exhibiting the lower Raman shift to the transition, and the peak at higher shift is identified as the process. To extract the ERS signal, we first subtract a linear background and normalize to a phonon peak, in this example (Fig. 3c) to the 2D + G peak at 4300 cm-1. After this we subtract the two spectra, which leaves us with the ERS signal. We fit the resulting ERS peaks by Gaussians to determine their Raman shift. The result of this procedure is shown in Fig. 3c. A detailed description of the fitting process can be found in section S6 of the Supplementary Information and the shared data 37.
We apply this procedure to samples with thicknesses varying from 3 to 12 graphene layers. Each sample undergoes meticulous examination using atomic force microscopy (AFM) to ensure accurate layer count. Additionally, we identify few layer samples through optical contrast measurements 32; 33. The ERS peaks of these samples can be seen in Fig. 4a. We can observe a clear trend: with increasing layer number the and peaks continuously shift to lower energy (lower Raman shift). By plotting these peak positions extracted by Gaussian fitting versus the layer number this trend is even more evident (see Fig. 4b). In certain cases the ERS peak positions have a larger estimated error (120 cm-1) due to lower signal to noise or due to the overlap of the ERS peak with the 2D mode in the case of 9 and 10 layers. These measurements align within our error margins with the calculated positions of ERS peaks from references: 27; 28 (see Fig. 4b). The measured ERS peak positions are also supplied as a table in the Methods (see Table 1).
@@
Next we check the homogeneity of the ERS peak position over the flake. Taking the difference in the crossed and parallel polarization measurements across the crystal, we plot the Raman intensity of the transition for the 8 and 7 layer areas. It becomes clear that the largest intensity of the ERS peak is located exactly on the rhombohedral 7 and 8 layer regions respectively (Fig. 5a). To check the homogeneity of the rhombohedral regions, in Fig. 5b we plot the position of the ERS peak. The standard deviation within both regions is under 5 cm-1. If a stacking fault were present in these regions, a change in peak energy two orders of magnitude greater would be expected. This is because the shift in ERS peak position, due to a stacking fault, is of the order of the change in energies with adding another layer to perfectly stacked RG 27. Thus, the change in peak position would be least 300 cm-1 in the presence of a stacking fault.
@@ Methods
We exfoliate samples using "blue tape" (Ultron systems P/N: 1008R-8.0), but other tapes work equally well. As substrate we used Si wafers (90 nm SiO2). Graphite samples were purchased from NGS Trading & Consulting GmbH (www.graphit.de). Based on hundreds of flakes investigated by Raman measurements, roughly 40% of them have some rhombohedral domains. We performed AFM measurements, using an NX10 microscope from Park Systems in tapping mode (non-contact mode).
@@
For Raman measurements we use a Witec 300rsa+ confocal Raman system, using 488, 532, and 633 nm laser excitations. All data shown in the main text are measured using a diffraction grating with 600 lines mm-1. Raman measurements are first analysed using the Witec data processing software supplied with the confocal Raman system. We performed final data processing and generated the figures, using the open-source Python tool: Ramantools 43.
| Layer Number | ERS peak position (cm-1) | ERS peak position (cm-1) |
@@
STM measurements are carried out at a temperature of 9 K, using an instrument from RHK (PanScan Freedom), with a base pressure of 510-11 Torr. Tunnelling conductance measurements were performed, using a Lock-in amplifier, at a frequency of 1372 Hz and a bias modulation amplitude of 5.5 mV. Samples investigated by STM are exfoliated flakes, supported on a Si/SiO2 substrate and contacted using In spikes 44. Data analysis of the STM measurements is carried out using the open-source Python tool: RHKPY 45.
@@
The optimized geometry and ground state Hamiltonian and overlap matrix elements of each structure were self consistently obtained by the SIESTA implementation of density functional theory (DFT) 46; 47; 48. SIESTA employs norm-conserving pseudopotentials to account for the core electrons and linear combination of atomic orbitals to construct the valence states. For all cases the considered samples were separated with a minimum of 1.35 nm thick vacuum in the perpendicular direction. The generalized gradient approximation of the exchange and the correlation functional was used with Perdew-Burke-Ernzerhof parametrization 49 with a double- polarized basis set. The geometry optimizations were performed until the forces were smaller than 0.1 eV nm-1 . The geometry of the considered structures were optimized for every configuration, initiated from the experimental in-plane lattice constant = 0.246 nm and out-of-plane lattice constant = 0.670 nm of hexagonal graphite. During the geometry relaxation the real-space grid was defined with an equivalent energy cutoff of 400 Ry and the Brillouin zone integration was sampled by a 1201201 Monkhorst-Pack -grid 50.